OpenStudy (anonymous):

A circle has its center at (-2, 5) and a radius of 4 units. What is the equation of the circle? x + 2)2 + (y - 5)2 = 16 (x + 2)2+ (y + 5)2= 16 (x + 2)2+(y - 5)2 = 4 (x - 2)2+(y + 5)2 = 4

OpenStudy (anonymous):

(x-h)2 + (y-k)2 = r2 gives you..? (h and k are the coordinates, (h,k))

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

replace (h,k) you have with the h and k in the equation.. dont forget minus signs

OpenStudy (anonymous):

The standard form for these circle equations is: (x-h)^2 + (y-k)^2 = r^2 You have h = -2, k = 5 and r = 4.

OpenStudy (anonymous):

PLug them in. Bam. Done.

OpenStudy (anonymous):

okay thanks